OpenStudy (anonymous):

WILL FAN & MEDAL!!!! please help! Determine which ordered pair is a solution of y= -5x+10. a. (-15, 5) b. (2, 0) c. (-1, -5) d. (3, 25)

OpenStudy (anonymous):

it is a because -15=-5(5)+10 -15=-15

rishavraj (rishavraj):

@jprose thts worng (-15 , 5) means x = -15 and y = 5

rishavraj (rishavraj):

hmmm for option A its not true.... now for option B (2 , 0) y = -5x + 10 0 = 0 so B is ur answer :))

OpenStudy (kitkat16):

There are actually three ordered pairs for this problem but B is your choice.